### 09 Apr 2026 v2

# How to Expose a Black Hole

Abstract
According to the correspondence principle of Horowitz and Polchinski, many black holes in string theory are continuously deformed to usual quantum systems involving D-branes and fundamental strings when the string coupling becomes sufficiently small. Therefore if we consider a configuration in space-time where the dilaton varies over an appropriate range, then a black hole moving in such a background will smoothly transition from the black hole state to a normal quantum state whose microstates are not hidden behind an event horizon. The possible obstruction to this mechanism comes from the fact that if the dilaton varies too fast then the adiabatic approximation may break down and / or the ambient space-time itself may collapse to a black hole and get hidden from the asymptotic observer. On the other hand, if the dilaton varies too slowly then the time that it takes for the black hole to travel the required distance will exceed the evaporation time of the black hole. We show that by choosing the background appropriately these obstructions can be avoided and a gentle motion towards the weak coupling region will convert the black hole into a normal quantum state without an event horizon.
1 Introduction
Our conventional picture of black hole formation and subsequent evaporation takes an asymmetric view: the black holes form by collapse of ordinary matter by classical process, but their evaporation happens by quantum process in the form of Hawking radiation, usually taking much longer than the time it takes for the formation. The goal of this paper will be to argue that in certain class of string theories we can partially avoid this asymmetry: A classical process can convert the black hole to a regular state without event horizon that can be probed by an external observer. Furthermore, the time needed to convert the black hole to a regular quantum state can be made parametrically small compared to what would be the evaporation time of the black hole if left unattended.
The main idea will be to use the correspondence principle formulated in [1, 2, 3], following earlier suggestions in [4, 5, 6, 7]. More recent perspective on this can be found in [8, 9, 10, 11]. This principle states that as we adiabatically change the string coupling from finite value to a sufficiently small value, a Schwarzschild black hole transforms to a highly excited elementary string state. This was generalized to a wide class of other charge carrying black holes, although at weak coupling these black holes correspond to systems of D-branes and fundamental strings instead of just fundamental strings.
Recently it was shown in [12, 13, 14] (see [15] for a different viewpoint) that in string theories with moduli fields in asymptotically flat space-time, one can produce finite energy background in which we have an arbitrarily large region in space-time inside which the moduli can take any desired values, in general different from their asymptotic values. More generally one can produce backgrounds in which the moduli vary from any desired value to any other, and the variation can be made arbitrarily slow.
This raises the possibility of realizing the black hole to elementary string / D-brane transition dynamically. Given a black hole, if the transition to the microscopic description happens around a value $g_{0}$ of the string coupling (which is a modulus of the theory) then by producing a background in which the coupling varies from a value much larger than $g_{0}$ to a value much smaller than $g_{0}$, and making the black hole roll in this background, we can achieve the desired transition. Once the black hole rolls to the weak coupling region, it is described by an ordinary quantum system. While determining the exact quantum state of the system may still be a challenging problem, it is no different from a high energy pure quantum state of matter that looks approximately thermal.
There are however some consistency conditions that need to be checked. While we can produce backgrounds in which the moduli vary arbitrarily slowly in space-time, there is an upper bound on how fast the moduli can vary. This comes from the fact that too fast a variation will make the background space-time collapse into a big black hole, and the whole system will get hidden behind an event horizon. Hence we have a lower limit on the time it takes for the black hole to travel from a region where the black hole description is valid to the region where the microscopic description is valid. Other lower bound on the transit time comes from the fact that in order that the adiabatic description is valid, the background must vary slowly in the frame of the black hole. At the same time, the time of travel must be small compared to the evaporation time of the black hole. We show that it is possible to satisfy all of these conditions together. Hence in such a background, a gentle roll towards the weak coupling region will convert the black hole to a regular quantum system whose microstates are not hidden behind an event horizon.
A similar mechanism for converting a black hole to an elementary string was considered in [11] with the help of large amplitude and large wavelength dilaton wave. However, as discussed in [14], there is an upper bound of order unity on the amplitude of a scalar wave in any dimensions, since for a wave of amplitude $a$ and wave-length $R$ measured in Planck units, the energy density is of order $(a/R)^{2}$ and the energy contained in a ball of size $R$ in $D$ space-time dimensions is of order $a^{2}R^{D-3}$. The associated Schwarzschild radius $a^{2/(D-3)}R$ will exceed $R$ for sufficiently large $a$, making the background collapse into a black hole. [^1] This is not a no go theorem since this argument does not take into account the full non-linear effects of gravity. However this shows that without having a more detailed analysis, we cannot trust the dilaton wave background to be horizon free. The background considered here does not suffer from this problem. Even though we use a large black hole to produce the desired background, the system that we study always remains outside the horizon of the large black hole and hence remains accessible to the asymptotic observer.
Note that the analysis of this paper is agnostic about the various solutions to the black information paradox[16, 17] that have been discussed in the literature. For example in the fuzzball proposal[18, 19] even at finite string coupling the horizon is replaced by horizonless geometry. The mechanism discussed in this paper would convert these horizonless geometries to regular weakly coupled quantum system made of strings and branes. In the island proposal[20, 21, 22, 23] for an old black hole the degrees of freedom inside the black hole are encoded in the outgoing Hawking radiation. Once the black hole rolls across the correspondence point, the horizon shrinks away and with that the island also disappears. In the holography of information proposal [24] one determines the quantum state of the black hole from the asymptotic region with the help of gravitational constraint equations. Just as we do not need to invoke this for finding the quantum state of an ordinary hot coal, similarly once the black hole crosses the correspondence point, one can study this just as we shall study an ordinary hot coal.
The rest of the paper is organized as follows. In section 2 we show how we can produce a background in which a Schwarzschild black hole can roll from the black hole state to a highly excited elementary string state. In section 3 we briefly discuss the case of other non-extremal black holes. In section 4 we discuss the case of BPS black holes. We end in section 5 with some comments.
2 Schwarzschild black hole
In this section we shall describe the dynamical transition of a Schwarzschild black hole to an elementary string excitation. In section 2.1 we review the correspondence principle for this system. In section 2.2 we describe the general strategy for producing the background that can be used to dynamically change a black hole to an elementary string state. In section 2.3 we describe explicit realization of such a background. In section 2.4 we describe an extension of the analysis in section 2.3 where the string coupling varies from a value of order unity all the way to a sufficiently small value where the black hole becomes an elementary string state.
2.1 Review of the correspondence principle
Let us work in a string compactification where we have $D$ non-compact space-time directions and let us denote by $g_{s}$ the string coupling in $D$ dimensions. We shall keep the size of the compact directions finite in string units and hence the string coupling in ten dimensions is also of order $g_{s}$. Let $\widetilde{m}$ be the mass of any state measured in string units and $m$ be the mass of the same state measured in $D$ dimensional Planck units. Newton’s gravitational constant is, by definition, of order unity in Planck units and is of order $g_{s}^{2}$ in string units. Then we have the relation:
$m\sim g_{s}^{2/(D-2)}\,\widetilde{m}\,$. (2.1)
Similarly, if $\widetilde{r}$ denote the size of the system measured in string units and $r$ denotes the same size measured in Planck units, then we have the relation
$r\sim g_{s}^{-2/(D-2)}\,\widetilde{r}\,$. (2.2)
In most of this section we shall use Planck units to express various quantities, but we can translate the results to string units using (2.1) and (2.2).
Let us consider an elementary string state at level $N$. For large $N$, the entropy of these states, defined as the logarithm of the degeneracy, varies as
$S\sim\sqrt{N}\,.$ (2.3)
The mass of the state in string units is given by
$\widetilde{m}\sim\sqrt{N}\,$, (2.4)
which translates to
$m\sim g_{s}^{2/(D-2)}\,\sqrt{N}$ (2.5)
in Planck units.
Let us now leave this system aside and consider a Schwarzschild or Kerr black hole of mass $m$ in Planck units. The radius $r_{s}$ of the black hole measured in Planck units is given by
$r_{s}\sim m^{1/(D-3)}\,.$ (2.6)
The entropy of such a black hole is of order
$S\sim r_{s}^{D-2}\sim m^{(D-2)/(D-3)}\,.$ (2.7)
The temperatures $T_{bh}$ of the black hole, measured in Planck units, is given by:
$T_{bh}\sim m^{-1/(D-3)}\,.$ (2.8)
The evaporation time $\tau_{bh}$ of the black hole, measured in the Planck units, is obtained by dividing the total mass by the product of the area of the horizon $\sim r_{s}^{D-2}$ and the energy flux $\sim T_{bh}^{D}$. This gives:
$\tau_{bh}\sim m^{(D-1)/(D-3)}\,.$ (2.9)
According to the correspondence principle[1, 2], at some particular value of $g_{s}$, the description of the system changes from that of a black hole to that of a fundamental string. The value of $g_{s}$ where it happens is determined by demanding that at the transition point the entropy and the mass of the black hole should match those of the elementary string states. Comparing (2.3), (2.5) with (2.7), we see that this gives
$\sqrt{N}\sim g_{s}^{2/(D-3)}N^{(D-2)/\{2(D-3)\}}\,$, (2.10)
and hence
$g_{s}\sim N^{-1/4}\,.$ (2.11)
Using (2.5), (2.6) and (2.11) we see that at this point
$m\sim N^{(D-3)/\{2(D-2)\}},\qquad r_{s}\sim N^{1/\{2(D-2)\}}\,.$ (2.12)
Using (2.2), (2.11) and (2.12) we now see that the radius of the black hole, measured in string units, is given by,
$\widetilde{r}_{s}\sim 1\,.$ (2.13)
Also from (2.8), (2.9), (2.11) and (2.12) we see that at the transition point
$T_{bh}\sim N^{-1/\{2(D-2)\}},\qquad\tau_{bh}\sim N^{(D-1)/\{2(D-2)\}}\,.$ (2.14)
Using (2.1) and (2.11), and using the fact that the conversion formula for the temperature is the same as that of the mass, we get the temperature of the black hole in string scale
$\widetilde{T}_{bh}\sim 1\,.$ (2.15)
From (2.13) and (2.15) we see that at the transition point the size of the horizon and the temperature of the black hole reaches the string scale, confirming that stringy effects become important at this point.
2.2 Dynamical transition from black hole to string
Let $\phi$ denote the dilaton field. We restrict our analysis to those string compactifications where $\phi$ is a modulus. Our goal will be to create a background in which $e^{\phi}$, that determines the local string coupling $g_{s}$, changes from $N^{-1/4}/\epsilon$ to $N^{-1/4}\epsilon$ for some small but fixed number $\epsilon$. The distance scale over which this happens should be much larger than the size of the black hole / string system we have analyzed in section 2.1, so that the black hole can be treated as a point particle moving in this background. Furthermore the various fields must vary sufficiently slowly so that the black hole moving through this background locally sees a configuration that is close to the vacuum. In particular the effect of any ambient radiation / other sources of stress tensor on the black hole should remain small, and the motion of the black hole can be taken to be adiabatic. Then as the black hole traverses this region, it will change into a fundamental string state whose internal state is not hidden from the asymptotic observer behind an event horizon. At the same time the size of the region over which the transition takes place cannot be arbitrarily large since the time taken for this process must be parametrically smaller than the evaporation time so that the black hole does not change appreciably during the transit. Furthermore, the energy of the required background should not be so high that it gets hidden behind an event horizon. Our goal will be to show that we can produce such a background.
We shall first describe the general strategy for achieving this and then illustrate this using a specific background. During this analysis we shall work with quantities measured in Planck units, but the final results of course will be independent of the units used in the analysis. Let us suppose that we have a (possibly time dependent) solution to the classical equations of motion in which we have a time-like trajectory along which $e^{\phi}$ changes from $N^{-1/4}/\epsilon$ to $N^{-1/4}\epsilon$, and the proper time along the trajectory does not scale with $N$ (when expressed in Planck units). This is a reasonable assumption since the net change in the canonically normalized field $\phi$ is independent of $N$. In section 2.3 we shall construct an explicit background satisfying this requirement and in section 2.4 we shall generalize this analysis to the case where $e^{\phi}$ changes from a finite value to $N^{-1/4}\epsilon$. We shall further assume that the beginning and the end points of the trajectory can send signals to the asymptotic observer. Then we can generate another solution to the classical equations of motion by scaling all covariant rank $k$ tensor fields by $\lambda^{k}$ and contravariant rank $k$ tensor fields by $\lambda^{-k}$, since under such a transformation a two derivative action scales by $\lambda^{D-2}$ in $D$ space-time dimensions and the equations of motion remain unchanged[25, 12, 13, 14]. In particular the $\sqrt{\det g}$ factor in the Lagrangian density produces a factor of $\lambda^{D}$ and the extra factor of $g^{\mu\nu}$ that contracts with the two derivatives produce a factor of $\lambda^{-2}$. Furthermore, the causal structure of space-time remains unchanged under this rescaling and the beginning and the end points of the trajectory can still send signals to the asymptotic observer. Since the metric scales by $\lambda^{2}$, the proper time taken during the travel scales as $\lambda$ and becomes large in the limit of large $\lambda$. In order to keep this small compared to the evaporation time $\tau_{bh}$ given in (2.14), we need
$\lambda<<N^{(D-1)/\{2(D-2)\}}\,.$ (2.16)
Note that on the right hand side of (2.16) we have used the evaporation time at the crossover point and not taken into account the effect of the small parameter $\epsilon$. This is because in the <<<< symbol it will be assumed implicitly that the inequality will hold even when the right hand side is multiplied by some appropriate power of $\epsilon$ to represent the worst case scenario.
(2.16) gives an upper bound on $\lambda$. We shall now derive a lower bound by requiring that the change in the state of the black hole is adiabatic during the transition. If the black hole moves at a speed of the order of the speed of light, then the distance covered per unit time is of the order of the radius of the black hole. We shall require that the background does not change appreciably over this distance. Since the distance over which the background changes scales as $\lambda$, this imposes the condition, using (2.12),
$\lambda>>r_{s}\sim N^{1/\{2(D-2)\}}\,.$ (2.17)
Under this condition the rate of change in the background fields, as seen from the perspective of the black hole, remains small and the transition can be called adiabatic. As a consequence of (2.17) the system also satisfies some other conditions that we shall now describe.
  1. 1.
    Since the length scale over which the background changes is large compared to the size of the black hole / string, we can describe the motion as that of a point particle moving in a background.
  2. 2.
    The acceleration of the black hole along the trajectory scales as $\lambda^{-1}$ and hence the Unruh temperature[26] seen by the black hole scales as $\lambda^{-1}$. [^2] Note that this is a crude estimate based on scaling and the actual acceleration and the associated temperature will depend on the details of the trajectory. While free fall along a geodesic provides a natural trajectory, this is not needed. For example the black hole may be gravitationally bound to another set of objects fitted with engines that slows down the fall. (2.17) ensures that this is small compared to the black hole temperature $N^{-1/\{2(D-2)\}}$ given in (2.14).
  3. 3.
    Since the local Unruh temperature is small compared to the black hole temperature, the total energy absorbed by the black hole during the transit is small compared to the total energy emitted by the black hole during the transit. The latter, in turn, is already small compared to the total mass of the black hole due to (2.16). Therefore the energy absorbed by the black hole during the transit remains small compared to its total mass.
  4. 4.
    A similar argument shows that the entropy absorbed by the black hole from the ambient radiation during the transit also remains small compared to its total entropy.
From this analysis we conclude that as long as (2.16) and (2.17) hold, the change in the state of the black hole can be taken to be adiabatic. Combining (2.16) and (2.17) we get,
$N^{1/\{2(D-2)\}}<<\lambda<<N^{(D-1)/\{2(D-2)\}}\,.$ (2.18)
This shows that as long as we can create a background in which the string coupling varies from $N^{-1/4}/\epsilon$ to $N^{-1/4}\epsilon$ for some small number $\epsilon$ that does not scale with $N$, we can scale the solution and by choosing the scaling parameter appropriately, we can ensure that the transition from black hole to elementary string is adiabatic and the total fractional change in the energy and entropy of the black hole during the transit remains small. The ability to produce such a varying dilaton configuration was discussed in [12, 13, 14] where it was shown that starting from any set of asymptotic values of the moduli, we can produce any other set of values inside an arbitrarily large region of space-time. Hence we can utilize the construction of [12, 13, 14] and then scale it by a parameter $\lambda$ satisfying (2.18) to produce a background that will change a black hole to an elementary string excitation. Once the system becomes an excitation of an elementary string, it can be probed by an external observer like any other quantum state.
Let us now discuss in what sense the state of the black hole does not change during the transit. Since the black hole radiates and absorbs radiation during the transit, the quantum state of the black hole itself changes during this process. To prevent this we would have to demand that the black hole does not emit or absorb even a single quantum of radiation during the transit, which will be too strong a condition to hold. For example if we demand that not even a single Hawking quanta is emitted by the black hole during the transit, or equivalently that the total entropy of the radiation emitted during the transit is small compared to unity, we get the condition
$\lambda\times T_{bh}^{D-1}\times r_{s}^{D-2}<<1\qquad\Rightarrow\qquad\lambda<<N^{1/\{2(D-2)\}}\,.$ (2.19)
This is in conflict with the lower bound on $\lambda$ given in (2.18). Indeed, since $r_{s}\sim N^{1/\{2(D-2)\}}$, this would imply that the transit time $\lambda$ is small compared to the size of the black hole which is clearly not compatible with the adiabaticity of the transition. However, some properties of the state remain unchanged during the transit. For example, if at the beginning of the roll the back hole was formed out of the collapse of a pure state, then at the end of the roll it will remain almost pure, in that its entanglement entropy will remain small compared to its thermodynamic entropy since due to the condition (2.16), the entropy emitted or absorbed by the black hole during the transit is parametrically smaller than $N^{1/2}$. On the other hand, if the black hole under consideration is an old black hole that is entangled with the Hawking radiation emitted in the past with an entanglement entropy of order $N^{1/2}$, then the entanglement entropy of the black hole does not change appreciably during the transit and at the end of the transit we shall get a regular quantum state of an elementary string without event horizon, entangled with the past Hawking radiation with an entropy of order $N^{1/2}$.
2.3 Varying dilaton background from charged black hole
We shall now describe an explicit example of a background of the type described above. This will be done with the help of a bigger black hole that we shall call the background black hole. For this we shall use the electrically charged black hole solution in heterotic or type IIA string theory reviewed in appendix A. The relevant part of the solution takes the form:
$\displaystyle ds^{2}$ $=$ $-(1+C\rho^{3-D})^{-2(D-3)/(D-2)}(1-2m\rho^{3-D})dt^{2}+(1+C\rho^{3-D})^{2/(D-2)}(1-2m\rho^{3-D})^{-1}d\rho^{2}$
$+$ $(1+C\rho^{3-D})^{2/(D-2)}\rho^{2}d\Omega_{D-2}^{2}\,$,
$e^{-2\phi}$ $=$ $(1+C\rho^{3-D})\,$,
$A_{t}$ $=$ $-{1\over\sqrt{2}}\,m\,\sinh\beta\,\rho^{3-D}\,(1+C\rho^{3-D})^{-1}\,.$ (2.20)
Here $\phi$ is the dilaton field, $A_{\mu}$ is a gauge field that couples to winding + momentum charge along a compact circle and $ds^{2}$ is the canonical Einstein frame metric. $m$ and $\beta$ are independent parameters, and
$C=m(\cosh\beta-1)\,.$ (2.21)
We shall take $\beta$ to be large so that $C>>m$ and work in the region
$m<<\rho^{D-3}<<C\,.$ (2.22)
In this region we have $m\sinh\beta\simeq C$ and the background black hole solution takes the form
$\displaystyle ds^{2}$ $\simeq$ $-(C\rho^{3-D})^{-2(D-3)/(D-2)}dt^{2}+(C\rho^{3-D})^{2/(D-2)}d\rho^{2}+(C\rho^{3-D})^{2/(D-2)}\rho^{2}d\Omega_{D-2}^{2}\,$,
$e^{-2\phi}$ $\simeq$ $(C\rho^{3-D})\,$,
$A_{t}$ $\simeq$ $-{1\over\sqrt{2}}\,(1-C^{-1}\rho^{D-3})\,.$ (2.23)
Note that in the expression for $A_{t}$ we have kept the leading non-constant term since the constant term gives vanishing field strength. Since the horizon of the black hole is at $\rho^{D-3}=2m$, all points in the range (2.22) are outside the horizon and remain causally connected to the asymptotic observer.
Now let us consider the effect of scaling by $\lambda$ described in section 2.2. Since the metric scales by $\lambda^{2}$, the gauge field scales by $\lambda$ and the dilaton remains constant, we see that we can achieve this by simply scaling $\rho$ and $t$ by $\lambda$ and $C$ by $\lambda^{D-3}$. Since the scaling of $\rho$ and $t$ are coordinate transformations, this means that we need to take $C$ to be large.
We can simplify the solution by introducing new coordinates $\bar{\rho}$ and $\bar{t}$ via
$\bar{\rho}=C\,\rho,\qquad\bar{t}=C^{3-D}\,t\,.$ (2.24)
In these coordinates the solution takes the form [^3] Since $F_{\bar{\rho}\bar{t}}$ has an extra factor of $C^{-1}$ and $C$ is large, one might wonder why we keep this term. The reason is that when we use the canonical metric, the gauge kinetic term is multiplied by a factor of $e^{-4\phi/(D-2)}$. This produces a factor of $C^{2}$ due to $C$ dependence of $\phi$ and cancels the $C^{-2}$ factor coming from the quadratic term in the gauge field.
$\displaystyle ds^{2}$ $\simeq$ $-\bar{\rho}^{\,2(D-3)^{2}/(D-2)}d\bar{t}^{2}+\bar{\rho}^{\,-2(D-3)/(D-2)}d\bar{\rho}^{\,2}+\bar{\rho}^{\,2/(D-2)}d\Omega_{D-2}^{2}\,$,
$e^{-2\phi}$ $\simeq$ $C^{D-2}\,\bar{\rho}^{\,3-D}\,$,
$F_{\bar{\rho}\bar{t}}$ $\simeq$ ${1\over\sqrt{2}}\,C^{-1}\,(D-3)\,\bar{\rho}^{\,D-4}\,.$ (2.25)
In this form the metric is $C$ independent and the scaling of the metric by $\lambda^{2}$ is generated by the transformation
$\bar{\rho}\to\lambda^{D-2}\bar{\rho},\qquad\bar{t}\to\lambda^{1-(D-3)^{2}}\bar{t}\,.$ (2.26)
In order to get $e^{\phi}\sim N^{-1/4}$, we need to take $\bar{\rho}\sim\bar{\rho}_{0}$ where
$C^{D-2}\,\bar{\rho}_{0}^{3-D}\sim N^{1/2}\,.$ (2.27)
In particular to have the coupling change from $N^{-1/4}/\epsilon$ to $N^{-1/4}\epsilon$ for some small number $\epsilon$, we need to have $\bar{\rho}$ change from $\bar{\rho}_{0}\epsilon^{-2/(D-3)}$ to $\bar{\rho}_{0}\epsilon^{2/(D-3)}$. In order to verify that $\rho\sim\bar{\rho}_{0}/C$ lies in the range (2.22), we note that according to (2.27), $\rho^{D-3}=(\bar{\rho}_{0}/C)^{D-3}\sim C\,N^{-1/2}$ and hence for large $N$ the upper bound $\rho^{D-3}<<C$ is satisfied automatically. On the other hand the lower bound $\rho^{D-3}>>m$ may be satisfied by requiring
$m<<C\,N^{-1/2}\,.$ (2.28)
Consider now the smaller black hole / string considered in section 2.1 and 2.2 rolling in this background black hole from $\bar{\rho}\sim\bar{\rho}_{0}\epsilon^{-2/(D-3)}$ to $\bar{\rho}\sim\bar{\rho}_{0}\epsilon^{2/(D-3)}$. In order to distinguish it from the background black hole, we shall call this the rolling black hole. Since the background black hole is much larger than the rolling black hole, we can treat the latter as a point mass moving in the field of the background black hole. During the transit the string coupling seen by the rolling black hole changes from $N^{-1/4}/\epsilon$ to $N^{-1/4}\epsilon$, as required for the black hole to string transition. It follows from (2.3) that the time taken for the roll, as measured by a local observer who is stationary with respect to the background, scales as $\bar{\rho}_{0}^{1/(D-2)}$ where we have ignored factors containing powers of $\epsilon$. Since the rolling black hole speed is not ultra-relativistic, the proper time measured by the rolling black hole also scales as $\bar{\rho}_{0}^{1/(D-2)}$. Requiring this to be smaller than the evaporation time $\tau_{bh}\sim N^{(D-1)/\{2(D-2)\}}$ given in (2.14), we get
$\bar{\rho}_{0}<<N^{(D-1)/2}\,.$ (2.29)
On the other hand, for the adiabatic approximation to hold, the time of transit $\bar{\rho}_{0}^{1/(D-2)}$ should be large compared to the size $r_{s}\sim N^{1/\{2(D-2)\}}$ of the horizon given in (2.12). This gives
$\bar{\rho}_{0}^{1/(D-2)}>>N^{1/\{2(D-2)\}}\qquad\Rightarrow\qquad\bar{\rho}_{0}>>N^{1/2}\,.$ (2.30)
Combining the two bounds, we get
$N^{1/2}<<\bar{\rho}_{0}<<N^{(D-1)/2}\,.$ (2.31)
(2.27) now gives
$N^{1/2}<<C<<N^{(D-2)/2}\,.$ (2.32)
This shows that in order to see the black hole to string transition for the rolling black hole, the background black hole parameter $C$ needs to be constrained.
Using these results we can also test the other conditions. For example using the lower bound on $C$ given in (2.32) we can verify that the mass of the background black hole, which is of order $C$, is much larger than the mass of the rolling black hole which is of order $N^{(D-3)/\{2(D-2)\}}$. Hence the backreaction of the rolling black hole on the background black hole geometry remains small. In particular, the causal structure of space-time remains unaffected and as long as the rolling black hole stays outside the horizon of the background black hole, given by the condition $\rho^{D-3}>2m$, it will be causally connected to the asymptotic observer. Also, it follows from (2.3) that for $\bar{\rho}\sim\bar{\rho}_{0}$, any scalar constructed from the fields and two derivatives scales as $\bar{\rho}_{0}^{-2/(D-2)}$. The only exception is $F_{\mu\nu}F^{\mu\nu}$, but the combination $e^{-4\phi/(D-2)}F_{\mu\nu}F^{\mu\nu}$ that appears in the Lagrangian density still scales as $\bar{\rho}_{0}^{-2/(D-2)}$. Hence for large $\bar{\rho}_{0}$, locally the rolling black hole experiences almost flat space-time. It also follows from this that a stationary observer in this background feels acceleration of order $\bar{\rho}_{0}^{-1/(D-2)}$. In the sense described in footnote 2, this also gives an order of magnitude estimate of the temperature experienced by the rolling black hole / string. From (2.31) we see that this is much smaller than the temperature $T_{bh}\sim N^{-1/\{2(D-2)\}}$ given in (2.14). Since (2.31) ensures that the effect of evaporation remains small during the transit, it now follows that the effect of absorption of the ambient radiation also remains small during the transit.
We can also check that the effect of Hawking radiation from the background black hole remains small. From (2.3) we can estimate the Hawking temperature $T_{BH}$ of the background black hole to be of order $C^{-1}m^{(D-4)/(D-3)}$. After taking into account the effect of the red-shift factor $(C\rho^{3-D})^{(D-3)/(D-2)}$ from (2.3), we get the red-shifted Hawking temperature at $\rho\sim\bar{\rho}_{0}/C$ to be
$T_{r}\sim T_{BH}\times(C\rho^{3-D})^{(D-3)/(D-2)}\sim C^{D-4}\,m^{(D-4)/(D-3)}\,\bar{\rho}_{0}^{-(D-3)^{2}/(D-2)}<<N^{-1/\{2(D-2)\}}\sim T_{bh}\,$, (2.33)
where in the last step we have used (2.27), (2.28) and (2.32). This shows that $T_{r}$ is small compared to the temperature $T_{bh}$ of the rolling black hole. Since the effect of the latter is small, we conclude that the effect of the absorption of Hawking radiation from the background black hole also remains small.
2.4 Fall from the asymptotic region
In sections 2.2 and 2.3 we have discussed how to make the black hole / string roll over a range where the effective string coupling $e^{\phi}$ varies from $N^{-1/4}/\epsilon$ to $N^{-1/4}\epsilon$. In this section we shall explore whether the rolling black hole can roll over a range of $e^{\phi}$ that starts at some small but $N$ independent value $\eta$ and then rolls all the way to the region $e^{\phi}\sim N^{-1/4}\epsilon$ in a time that is small compared to the decay time of the black hole. For this we have to start at a value $\bar{\rho}_{1}$ of $\bar{\rho}$ where $e^{\phi}\sim\eta$. (2.3) gives
$C^{D-2}\bar{\rho}_{1}^{3-D}\sim\eta^{-2}\,.$ (2.34)
Comparing this with (2.27) we get,
$\bar{\rho}_{1}\sim\bar{\rho}_{0}\,(N\eta^{4})^{1/\{2(D-3)\}}\,.$ (2.35)
Note that $\bar{\rho}_{0}$ is still defined to be the value of $\bar{\rho}$ where $e^{\phi}\sim N^{-1/4}$. Using (2.3) we see that the total transit time in the frame of the rolling black hole from $\bar{\rho}\sim\bar{\rho}_{1}$ to $\bar{\rho}_{0}$ is of order
$\bar{\rho}_{1}^{1/(D-2)}\sim\bar{\rho}_{0}^{1/(D-2)}\,(N\eta^{4})^{1/\{2(D-3)(D-2)\}}\,$, (2.36)
with most of the time spent in the region $\bar{\rho}\sim\bar{\rho}_{1}$. Requiring this to be small compared to the evaporation time $\tau_{bh}\sim N^{(D-1)/\{2(D-2)\}}$ gives
$\bar{\rho}_{0}<<\eta^{-2/(D-3)}N^{(D-1)/2-1/\{2(D-3)\}}\,.$ (2.37)
The lower bound on $\bar{\rho}_{0}$ that arises from requiring that the rolling is adiabatic also gets modified. For this we note from (2.3) that between $\bar{\rho}_{1}$ and $\bar{\rho}_{0}$ there is a redshift factor
$\gamma\sim(\bar{\rho}_{1}/\bar{\rho}_{0})^{(D-3)^{2}/(D-2)}\sim(N\eta^{4})^{(D-3)/\{2(D-2)\}}\,$, (2.38)
where we used (2.35). Hence a freely falling rolling black hole will have most of its mass converted to kinetic energy by the time it falls to the position $\bar{\rho}_{0}$ and it will move with ultra-relativistic speed. Hence at $\bar{\rho}\sim\bar{\rho}_{0}$, the rolling black hole has time dilation by a factor of $\gamma$ and the proper time of transit from $e^{\phi}\sim N^{-1/4}/\epsilon$ to $e^{\phi}\sim N^{-1/4}\epsilon$ is given by $\bar{\rho}_{0}^{1/(D-2)}/\gamma$ instead of $\bar{\rho}_{0}^{1/(D-2)}$. For the adiabatic approximation to be valid, we need this to be small compared to the size $r_{s}$ of the rolling black hole. This changes (2.30) to
$\bar{\rho}_{0}^{1/(D-2)}(N\eta^{4})^{-(D-3)/\{2(D-2)\}}>>N^{1/\{2(D-2)\}}\,.$ (2.39)
This gives
$\bar{\rho}_{0}>>N^{(D-2)/2}\,\eta^{2(D-3)}\,.$ (2.40)
There are various other equivalent ways of arriving at the same conclusion, e.g. in the rest frame of the rolling black hole, the ambient geometry will have a length contraction by a factor of $\gamma$, and we need to demand that even after taking this into account, the fields in the ambient space-time should vary over a length scale much larger than the size of the rolling black hole.
Combining (2.40) with (2.37) we get
$N^{(D-2)/2}\eta^{2(D-3)}<<\bar{\rho}_{0}<<\eta^{-2/(D-3)}N^{(D-1)/2-1/\{2(D-3)\}}\,.$ (2.41)
For $D\geq 5$ and large $N$, the upper bound is larger than the lower bound, and hence we can find $\bar{\rho}_{0}$ satisfying these conditions. For $D=4$, the $N$ dependence of the two sides are identical, but by taking $\eta$ to be small, we can still ensure that the upper bound is larger than the lower bound. Thus we see that it is possible for the rolling black hole to roll all the way from a region where the string coupling is small but $N$ independent to the region where it becomes an elementary string excitation. On the other hand, if the rolling black hole undergoes a controlled fall as discussed in footnote 2, maintaining its speed at order unity but not parametrically close to unity, then the lower bound on $\bar{\rho}_{0}$ will be given by the earlier bound $N^{1/2}$, and even for $D=4$ and finite $\eta$, the upper bound on $\bar{\rho}_{0}$ remains larger than the lower bound for large $N$.
Finally we can consider the fall from the asymptotic region $\rho\sim C^{1/(D-3)}$. In this case the transit time is of order $C^{1/(D-3)}$. Comparing this with the transit time $\bar{\rho}_{1}^{1/(D-2)}\sim C^{1/(D-3)}\eta^{2/\{(D-2)(D-3)\}}$ computed from (2.36) and (2.34), we see that this corresponds to setting $\eta\sim 1$ in the previous analysis. Also from (2.3) and (2.27) we see that the redshift factor $\gamma$ at a position $\rho\sim\bar{\rho}_{0}/C$ is now of order $(C\rho^{3-D})^{(D-3)/(D-2)}\sim N^{(D-3)/\{2(D-2)\}}$. Comparing this with (2.38) we again see that that this corresponds to setting $\eta\sim 1$ in the previous analysis. Thus for $D\geq 5$, even a free fall of the rolling black hole from the asymptotic region will convert a black hole to an elementary string state in a controlled fashion. For $D=4$ the adiabatic approximation will break down near the end of the roll, but this can be avoided by arresting the fall by external force as described in footnote 2 so that the rolling black hole does not become ultra-relativistic during the roll. In the extreme case, if we bring the rolling black hole at rest at $\bar{\rho}\sim\bar{\rho}_{0}\epsilon^{2/(D-3)}$ where it becomes an elementary string state, the proper acceleration felt by the rolling black hole will be of order $\bar{\rho}_{0}^{-1/(D-2)}\epsilon^{-2/\{(D-2)(D-3)\}}$. This remains small compared to the string scale as long as (2.30) holds. An observer in the rest frame of this black hole can study its quantum state and send the information back to the asymptotic observer since both the rolling black hole and the observer will be outside the global event horizon (which in the approximation we are using, is the horizon of the background black hole at $\rho^{D-3}=2m$).
3 Other non-extremal black holes
In this section we shall briefly discuss the case where the rolling black hole is a non-extremal black hole carrying charge(s) and / or angular momenta. We shall assume that the black hole is finite distance away from extremality. In that case the scaling of temperature, entropy and other thermodynamic quantities as a function of mass have the same form as in the case of a Schwarzschild black hole, and the analysis follows a path that is more or less identical to that in section 2, except that the large parameter $N$ has to be interpreted as the square of the entropy $S$ so that (2.3) holds:
$S\sim\sqrt{N}\,.$ (3.1)
If $r_{s}$ denotes the horizon radius measured in the Planck scale, then we have
$r_{s}\sim S^{1/(D-2)}\sim N^{1/\{2(D-2)\}}\qquad\Rightarrow\qquad\widetilde{r}_{s}\sim g_{s}^{2/(D-2)}N^{1/\{2(D-2)\}}\,$, (3.2)
where we used (2.2) in the second step. On the other hand, it was shown in [1] that at the correspondence point, the size of the horizon, measured in string scale, is of order unity, and hence (2.13) holds:
$\widetilde{r}_{s}\sim 1\,.$ (3.3)
Comparing (3.2) and (3.3), we get
$g_{s}\sim N^{-1/4}\,$, (3.4)
at the correspondence point. This is the same relation as (2.11). The rest of the analysis now follows as in section 2. The only other difference is that the rolling black hole is also a source of the dilaton and other scalars and the gauge fields, but for black holes that are not near extremal, the changes in various equations are of order unity and can be ignored.
For near extremal black holes the situation is more complicated since the temperature does not scale as (2.8) and we have a new large parameter. Since the effect of this is to increase the evaporation time of the black hole, we expect that the upper bound on the transit time will be larger and it will be easier to construct background in which the system transforms from a black hole state to the regular quantum state. However, as pointed out in [1], for some near extremal systems the correspondence principle has the puzzling feature that as we reduce the string coupling, even before reaching the correspondence point the black hole description breaks down since the curvature of the string metric reaches the string scale in some region between the horizon and the asymptotic region. For this reason we shall not study these cases in detail. Instead in the next section we shall focus on the BPS black holes.
4 BPS black holes
BPS black holes have been used to count microstates of black holes to high precision. Since appropriate supersymmetric index of the black hole is expected to be protected under a change in the coupling constant, we can carry out the counting at weak coupling where the system is described as an usual quantum state containing D-branes, fundamental strings and other known objects, while at strong coupling the system is described as a black hole[27]. Hence by exploiting the varying dilaton background described in section 2 we can make the transition between the black hole and microscopic description dynamical. As the black hole rolls from the strong to the weak coupling region, its description changes from that of a regular extremal BPS black hole to a system containing D-branes and other objects with regular quantum mechanical description.
In fact, for BPS black holes the situation is better than that for non-BPS black holes discussed in the previous sections since the BPS black holes have zero temperature and do not evaporate. Hence there is no upper limit on the time of rolling. By taking the scaling parameter $\lambda$ of section 2.2 (or equivalently the parameter $C$ of the background black hole of section 2.3) to be sufficiently large one can ensure that a not a single quantum of ambient Hawking or Unruh radiation is absorbed by the rolling black hole and hence the state of the rolling black hole remains unchanged during the motion. This can be seen by noting that the ambient temperature of the background scales as $\lambda^{-1}$ and hence the ambient entropy density scales as $\lambda^{-(D-1)}$. This has to be multiplied by the travel time $\lambda$ and the area of the rolling black hole to compute the total entropy absorbed during the transit. Since the area of the rolling black hole does not scale with $\lambda$, we see that the net entropy absorbed during the transit scales as $\lambda^{-(D-2)}$ and can be made as small as we like by taking $\lambda$ large. In particular when it becomes much smaller than one, it will imply that not even a single quantum of ambient radiation is absorbed by the black hole during its motion.
[…]
5 Discussion
In our analysis we have shown that the black hole can be transformed into an ordinary quantum system containing stings and branes by letting it drop to a region close to a much bigger near extremal electrically charged black hole. We have not discussed how to extract information from this system, but since this is an ordinary quantum system, an observer falling with the rolling black hole should be able to study the quantum state of this system using standard methods and transmit the data to an asymptotic observer. The time scale of such experiments will be long due to the weakness of the string coupling and one might wonder whether the rolling black hole will fall into the background black hole before such experiments could be performed. Furthermore as the coupling becomes weaker as we approach the horizon of the background black hole the time scale of measurement becomes longer. The latter problem could be avoided by adding a small amount of magnetic charge to the background black hole so that the attractor value of the string coupling is given by $N^{-1/4}/\epsilon$ for some fixed small number $\epsilon$. In that case, even if the rolling black hole falls towards the horizon of the background black hole, the coupling does not decrease any further. Also one could in principle arrest the fall of the rolling black hole towards the background black hole using the classical gravitational field of other (accelerating) objects placed in appropriate regions of space-time (see footnote 2 for a discussion on this).. While we have not studied this problem in detail, we expect that this should be possible in principle as long as the system of which the rolling black hole is a part remains outside the horizon of the background black hole.
Since black holes are expected to have chaotic spectrum while the spectrum of weakly coupled strings and branes have regular pattern, one might wonder how the transition from black hole to the brane description affects the spectrum. To this end, note that the spectrum of any microscopic system made of branes and strings is renormalized by quantum corrections. If the quantum corrections are chaotic, this could turn a regular spectrum into a chaotic spectrum. In particular it was shown in [28] that the interaction between the string states is highly chaotic. Since the renormalized mass depends on the interaction, it is quite plausible that the renormalized masses will display a chaotic spectrum.
Acknowledgement: I wish to thank Nejc Ceplak, Roberto Emparan, Andrea Puhm, Marija Tomasevic and Yoav Zigdon for useful communications. This work was supported by the ICTS-Infosys Madhava Chair Professorship and the Department of Atomic Energy, Government of India, under project no. RTI4019.
Appendix A Electrically charged black hole solution
In this appendix we shall review the construction of the electrically charged, non-rotating black hole solution of string theory in $D$ non-compact space-time dimensions. First we shall consider heterotic string theory compactified on a six dimensional torus $T^{6}$, and then show how the same construction can be generalized to any string compactification for which the compact space has a circle factor. The massless fields of this theory consist of the canonical metric $ds^{2}$, the two form field, the four dimensional dilaton $\phi$, a set of scalars taking values in $O(6,22)/O(6)\times O(22)$, encoded in a symmetric $O(6,22)$ matrix $M$ and 28 gauge fields, collectively denoted by a 28 dimensional vector $A_{\mu}$. We denote by $\widetilde{ds}^{2}=e^{2\phi}ds^{2}$ the string metric. The background black hole will be taken to be an electrically charged non-rotating black hole solution, given by[29]
$\displaystyle\widetilde{ds}^{2}$ $=$ $-\Delta^{-1}\,\rho^{2}(\rho^{2}-2m\rho)dt^{2}+\rho^{2}\,(\rho^{2}-2m\rho)^{-1}d\rho^{2}+\rho^{2}\,d\Omega_{2}^{2}\,$,
$e^{-2\phi}$ $=$ $\Delta^{1/2}\,\rho^{-2}\,$,
[…]
where
$\displaystyle\Delta$ $\equiv$ $\rho^{4}+2m\rho^{3}(\cosh\alpha\cosh\beta-1)+m^{2}\rho^{2}(\cosh\alpha-\cosh\beta)^{2}\,$,
$P$ $\equiv$ $2\,\Delta^{-1}\,m^{2}\rho^{2}\sinh^{2}\alpha\sinh^{2}\beta\,$,
$Q$ $\equiv$ $-2\Delta^{-1}m\rho\sinh\alpha\sinh\beta\{\rho^{2}+m\rho(\cosh\alpha\cosh\beta-1)\}\,$, (A.2)
$\vec{n}$ is a 22 dimensional unit vector and $\vec{p}$ is a six dimensional unit vector. The 2-form field vanishes. $m$, $\alpha$ and $\beta$ are arbitrary constants, that, together with $\vec{n}$ and $\vec{p}$, label the mass and the charges carried by this background black hole.
We set
$\alpha=0,\qquad m(\cosh\beta-1)=C\,$, (A.3)
so that we have
$\Delta=\rho^{4}\left(1+C\rho^{-1}\right)^{2},\qquad P=0,\qquad Q=0\,.$ (A.4)
Then (A) takes the form:
$\displaystyle\widetilde{ds}^{2}$ $=$ $-(1+C\rho^{-1})^{-2}(1-2m\rho^{-1})\,dt^{2}+(1-2m\rho^{-1})^{-1}d\rho^{2}+\rho^{2}d\Omega_{2}^{2}\,$,
$e^{-2\phi}$ $=$ $(1+C\rho^{-1})\,$,
$M$ $=$ $I_{28}\,$,
[…]
Physically $M$ denotes the moduli associated with the components of the metric, the two form field and the ten dimensional gauge fields along $T^{6}$ and $A_{\mu}$ denotes the four dimensional gauge fields associated with the components of the metric and two form field with one index along internal direction and one index along the non-compact direction and the components of the ten dimensional gauge field along the non-compact directions. In particular the solution described above carries equal amount of fundamental string winding and momentum charges along the direction $\vec{p}$ in $T^{6}$. [^4] The equality of momentum and winding charges is not necessary for our analysis. For any other finite ratio of these charges, $M$ would approach some other constant matrix at small $\rho$.
From this description it is clear that the same solution can be lifted to any compactification to 3+1 dimensions in which we have one circle direction among the compact coordinates. $M$ is in any case frozen to a constant and we simply have to interpret $A_{t}$ as the gauge field that couples to both momentum and winding charge along the circle. To lift it to a compactification with $D$ non-compact space-time dimensions with at least one circle among the compact coordinates, we need to recall how the solution was constructed in [29]. The essential idea was to start with a Schwarzschild black hole in four space-time dimensions and then perform a duality rotation that mixed the time direction with the compact directions[30, 31, 32]. The same operation can be carried out in $D$ space-time dimensions by starting with a Schwarzschild black hole in $D$ space-time dimensions. The only difference will be that in the initial solution all factors of $m/\rho$ would be replaced by $m/\rho^{D-3}$ and $d\Omega_{2}^{2}$ will be replaced by $d\Omega_{D-2}^{2}$. This will give the solution for $D$ non-compact space-time directions[33]:
$\displaystyle\widetilde{ds}^{2}$ $=$ $-(1+C\rho^{3-D})^{-2}(1-2m\rho^{3-D})\,dt^{2}+(1-2m\rho^{3-D})^{-1}d\rho^{2}+\rho^{2}d\Omega_{D-2}^{2}\,$,
$e^{-2\phi}$ $=$ $(1+C\rho^{3-D})\,$,
$A_{t}$ $=$ $-{1\over\sqrt{2}}\,m\,\sinh\beta\,\rho^{3-D}\,(1+C\rho^{3-D})^{-1}\,.$ (A.6)
We have dropped the expression for $M$ since this is a constant anyway, and also not displayed the internal vector $\vec{p}$ associated with the gauge field, with the understanding that we are considering a solution carrying equal amount of momentum and winding charges along one of the compact circles. The canonical $D$ dimensional metric $ds^{2}=e^{-4\phi/(D-2)}\widetilde{ds}^{2}$ now takes the form:
$\displaystyle ds^{2}$ $=$ $-(1+C\rho^{3-D})^{-2(D-3)/(D-2)}(1-2m\rho^{3-D})dt^{2}+(1+C\rho^{3-D})^{2/(D-2)}(1-2m\rho^{3-D})^{-1}d\rho^{2}$ (A.7)
$+$ $(1+C\rho^{3-D})^{2/(D-2)}\,\rho^{2}\,d\Omega_{D-2}^{2}\,.$